Grand Palace Hotel demand function for its guest room for the month of May 2021 is represented by the following equation.

Qd=700–2Pr +0.21TS+1.6I+0.05E Where,

Grand Palace Hotel and traveller’s expectation that a lower room price over the next six months:

TS = 20

I = 2.5

E = 100

  1. Derive the demand equation for Grand Palace Hotel for the month of May 2021, based on the assumption of ‘ceteris paribus’.
  2. Calculate the price for a room when the demand is 150
  3. Defend whether the assumption that the principle of ‘ceteris paribus’

is a reasonable assumption.


1
Expert's answer
2021-04-14T11:23:34-0400

"Qd=700-2Pr+0.21TS+1.6I+0.05"


"Qd= 700-2Pr+4.2+4+5"


"Qd=713.2-2Pr"


"150=713.2-2Pr"


Like terms:


"2Pr=713.2-150"


"2Pr=563.2"


Price=281.6


3.) The principle ceteris peribus is only valid when other factors are held constant. Otherwise, it is operate in reality.